HIGH-RESOLUTION SPECTROSCOPY OF np RYDBERG STATES OF He_{2}: 1. RYDBERG-STATE-RESOLVED THRESHOLD IONIZATION SPECTRA OF METASTABLE He_{2}

\underline{JINJUN LIU}, DANIEL SPRECHER, MATTHIAS RAUNHARDT, MARTIN SCHÄFER, and FRÉDÉRIC MERKT, ETH Zürich, Laboratorium für Physikalische Chemie, Wolfgang-Pauli-Strasse 10, 8093 Zürich, Switzerland.

A supersonic beam of metastable He_{2}^{∗} a ^{3}Σ_{u}^{+} molecules was generated using a pulsed discharge at the exit of a pulsed valve prior to the gas expansion into vacuum.^{a} Transitions to high np Rydberg states were recorded using photoionization and Rydberg-state-resolved threshold ionization spectroscopy.^{b} Overview scans at moderate resolution (0.3 cm^{−1}) were obtained with ionization fields ranging from 1.3 to 133 V/cm, lowering the ionization thresholds by 5.5 and 55 cm^{−1}, respectively. Using a solid-state UV laser system^{c} with a 20 MHz bandwidth, high-resolution spectra of Rydberg series with n up to 150 and with resolved fine structure of the initial He_{2}^{∗} a ^{3}Σ_{u}^{+} (N′′) state were recorded. The assignment of the observed Rydberg states is based on multichannel quantum defect theory calculations from a recent study^{a} of pulsed-field-ionization zero-kinetic-energy (PFI-ZEKE) photoelectron and photoionization spectra of He_{2} (see following talk). The extrapolation of the observed Rydberg series to their limits enabled the determination of the ionization energy of the a ^{3}Σ_{u}^{+} state and the rotational structure of the He_{2}^{+} ion with a precision of better than 20 MHz.

aM. Raunhardt, M. Schäfer, N. Vanhaecke, F. Merkt, J. Chem. Phys. 128, 164310 (2008).

bR. Seiler, U. Hollenstein, G. M. Greetham, F. Merkt, Chem. Phys. Lett. 346, 201 (2001).

cR. Seiler, Th. A. Paul, M. Andrist, F. Merkt, Rev. Sci. Instr. 76, 103103 (2005).